OpenStudy (anonymous):

Enter the value of the expression. -1.2*(2+6)=-9.6 right?

OpenStudy (anonymous):

@mathmale i have a math question

OpenStudy (anonymous):

@naga447

OpenStudy (mathmale):

-1.2*(2+6)=-9.6? Due to Order of Operations rules, must evaluate the quantity inside parentheses first. What is 2+6? Throw out the 2+6 and replace it, inside the parentheses, by your sum.

OpenStudy (anonymous):

ok so i have to do 2+6 first cause of oder of operation

OpenStudy (mathmale):

Yes, because 2+6 shows up INSIDE PARENTHESES. Your answer?

OpenStudy (anonymous):

its 8 2+6 is 8 but now i have to times it by -1.2 right

OpenStudy (mathmale):

2+6 = 8, yes. So, -1.2*(2+6)=-9.6 becomes -1.2*(8), and yes, you do need to multiply now.

OpenStudy (anonymous):

ok so te answer is -9.6

OpenStudy (mathmale):

Yes. Review and remember Order of Operations rules.

OpenStudy (mathmale):

some people remember them as PEMDAS.

OpenStudy (anonymous):

thank u